OpenStudy (anonymous):

If (a+4b)/(5a-7b)=2, find the value of (3a-7b)/(a+b).

OpenStudy (anonymous):

(a+4b)/(5a-7b)=2 a+4b=10a-14b 9a=18b a=2b

OpenStudy (anonymous):

right?

OpenStudy (zzr0ck3r):

solve for a, then use that to solve for the other expression

OpenStudy (anonymous):

Now, substitute a=2b in (3a-7b)/(a+b)

OpenStudy (anonymous):

CAn u?

OpenStudy (anonymous):

@sauravshakya I think I would end with variables in the answer. My teacher told me it would be a number value.

OpenStudy (anonymous):

Yes the answer will be number value

OpenStudy (anonymous):

@zzr0ck3r Can you explain how?

OpenStudy (anonymous):

Just try it

OpenStudy (zzr0ck3r):

just work it through

OpenStudy (anonymous):

@sauravshakya OH! I see now. Thank you!!

OpenStudy (anonymous):

Welcomx